OpenStudy (anonymous):

The sum of the digits in a two digit number is 9. If the digits are reversed, the new number will be 9 less than the original number. What is the original number?

OpenStudy (anonymous):

let the digits be x and y x+y= 9 x=9-y hence the number would be 10(9-y)+y is the number now if the digits are reversed we got 10y+ (9-y) since y would be tens digit hence we would multiply it with 10 now see it is given that 10(9-y) + y - (10y + 9-y) 90-10y + y - 9y - 9 81 -18y = 9 -18y = 9 -81 -18y = -72 y = -72/-18 y=4 so the digits are 9-y and y one is 4 and the other one is (9-4) = 5 the original number was 45 understood if any other problem tell me

OpenStudy (anonymous):

also, 12 and 21 work

OpenStudy (anonymous):

yes alex

OpenStudy (anonymous):

i find the easiest method for these is to brute logic them. but it is important to learn the equations and understand the principles behind them

OpenStudy (anonymous):

yes u r right alex

OpenStudy (anonymous):

alex how does 12 and 21 work?? sum of digits must be 9...

OpenStudy (anonymous):

yes KBOY is also right alex 12 and 21 is not the posssible answer